OpenStudy (anonymous):

Simplify (3b2 – 2b + 3) – (5b2 + 4b – 2)

OpenStudy (anonymous):

8b2 – 6b + 5 8b2 – 2b + 5 –2b2 – 6b + 5 –2b2 – 2b + 5

jimthompson5910 (jim_thompson5910):

(3b^2 – 2b + 3) – (5b^2 + 4b – 2) 3b^2 – 2b + 3 – 5b^2 - 4b + 2 (3b^2– 5b^2)+(-2b-4b)+(3 + 2) -2b^2-6b+5 So the answer is -2b^2-6b+5

OpenStudy (anonymous):

Okaye thanks

OpenStudy (angela210793):

I have forgotten how to add :'(

OpenStudy (anonymous):

What is the slope of the line which passes through (2, 6) and (0, 2)?

jimthompson5910 (jim_thompson5910):

m = (y2-y1)/(x2-x1) m = ( 2- 6)/(0-2) m = (-4)/(-2) m = 2 So the slope of the line that passes through the two points is 2

OpenStudy (anonymous):

2 –3 4 –2

OpenStudy (anonymous):

What is the degree of the polynomial 3h2 – 4h + 1?